Ray SU bisects angle RST. The measure of angle RSU=3x-1, the measure of angle UST=2x+13 and the measure of angle RST=6x-2. Find x.
Answer by cleomenius(959)

Since SU bisects RST, angle RSU plus angle UST equals angle RST.
We are given RST = 6x -2
We are given RSU = 3x -1
We are given UST = 2x + 13
Therefore, 3x - 1 + 2x + 13 = 6x -2
5x + 12 = 6x - 2
14 = x
RST = 6(14) - 2 = 82
RSU = 3(14) - 1 = 41
UST = 2x + 13 = 41
Therefore, this does check.
Cleomenius